OpenStudy (anonymous):

Events A and B are independent. P(A)=0.4 and P(B)=0.3. Find P(A∪B) to two decimal places.

OpenStudy (anonymous):

since A and B are independent, \[P(A \cup B) = P(A) + P(B)\]

OpenStudy (anonymous):

i thought it was P(a) + P(b) - ((P(a)*P(b))

OpenStudy (anonymous):

sorry my bad, that's if they are mutually exclusize

OpenStudy (anonymous):

yes, since A and B are independent \[P(A \cap B) = P(A) * P(B)\]

OpenStudy (kropot72):

\[P(A \cup B)=0.4+0.3-(0.4\times 0.3)\]

OpenStudy (deoxna):

@kropot72 but how do you know they are combined events?

OpenStudy (kropot72):

If A and B are two events, then A union B is the event A or B (or both). The probability of A or B is given by: \[P(A \cup B)=P(A)+P(B)-P(A\ intersection\ B)\]

OpenStudy (deoxna):

But how do you know they are not mutually exclusive? Events can be both independent and mutually exclusive... (sorry if I'm being a pain but I really do want to get this)

OpenStudy (kropot72):

Events A and B are independent if, and only if, \[P(A\ intersection\ B)=P(A)\times P(B)\] If A and B are mutually exclusive events the probability of A and B occurring is zero, \[P(A\ intersection\ B)=0\] In this question it is given that events A and B are independent.

OpenStudy (deoxna):

Wow, I really need to brush up on my probability... Ok thanks!